OpenStudy (anonymous):

Frank needs to fill juice cups for his little sister's birthday party. There are 35 juice cups and each juice cup holds 12 fluid ounces. The juice comes in 1-gallon bottles. How many 1-gallon bottles of juice will Frank need to purchase? (1 quart = 32 fluid ounces, 1 gallon = 4 quarts) 1 bottle 2 bottles 3 bottles 4 bottles

OpenStudy (mrnood):

First work out how many fluid ounces are required 1 cup = 1 * 12 fl oz 2 cup = 2 * 12 fl oz = 24 fl oz 3 cup = 3 * 12 fl oz = 36 fl oz . . 35 cup = ? * 12 fl oz = ?? fl oz post that step here an I will help you get the rest of the answer

OpenStudy (anonymous):

im sooo confused

OpenStudy (anonymous):

35 cup=35*12

OpenStudy (anonymous):

=420 fl oz

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

4 bottles

OpenStudy (anonymous):

First step: figure out how many fluid ounces in a gallon. 1 gallon=4 quarts=132 fluid ounces How many fluid ounces total do you need? 35 cups multiplied by 12 fluid ounces is 420 fluid ounces. How many 132 fluid ounce jugs do you need to get 420 fluid ounces filled? If you need 420 and you buy bottles that contain 132... how many times does 132 go into 420? so... 420/132= 3.1818182 round that up and you get (4)

OpenStudy (anonymous):

do you get it @Fallingstar80

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome!